OpenStudy (y2o2):

A body of mas 20 kg. is suspended from the end of a spring balance fixed in a balloon.find the apparent weight of the body indicated in the balance in kg.wt. in each of the following cases: a) the balloon moves upwards with a uniform velocity 2.5 m/s . b) the balloon moves upwards with a uniform acceleration 4.9 m/s² . c) the balloon moves upwards with a uniform acceleration 4.9 m/s². (g = 9.8 m/s²)

OpenStudy (anonymous):

\[\sum F = ma\] here, \[\sum F =F_s-mg\] You want \(F_s\)

OpenStudy (anonymous):

 (a)if baloon moves upward with uniform velocity then there is no external acceleration present ,that is no pseudo force here, simply ,apparent weight of the body(k*x)=mg=20*9.8=196N (b)if there is an external acceleration in the upward direction i.e in the direction of the motion of the balloon therefore the direction of pseudo force will be in downward : hence , apparent weight of the body(kx)=mg+ma(net downward force)=196+20*4.9=294N (c)if there is an external acceleration in the downward direction i.e in the opposite direction of the motion of the balloon therefore the direction of pseudo force will be in upward : hence , apparent weight of the body(kx)=mg-ma(net downward force)=196-20*4.9=98N if you have any query,ask me.......